2. Please show all of your work when answering the following:

I. Use Chebyshev’s theorem to find what percent of the values will fall between 252 and 336 for a data set with a mean of 294 and standard deviation of 14.

II. Use the Empirical Rule to find what two values 95% of the data will fall between for a data set with a mean of 223 and standard deviation of 11

Answer by stanbon(75887)

Please show all of your work when answering the following:
I. Use Chebyshev’s theorem to find what percent of the values will fall between 252 and 336 for a data set with a mean of 294 and standard deviation of 14.
---
(252-294)/14 = -3
(336-294)/15 = +3
----
According to Cheb, (1-(1/3)^2)% = 89% of the data lies within 3 standard
deviations of the mean.
--------------------------
II. Use the Empirical Rule to find what two values 95% of the data will fall between for a data set with a mean of 223 and standard deviation of 11
--
95% of the data lies within 2 standard deviations of the mean.
Lower limit: 223-2*11 = 201
Upper limit: 223+2*11 = 245
==================================
Cheers,
Stan H.